# Quantum Information meets High-Energy Physics: Input to the update of the European Strategy for Particle Physics

### v1

@@
Abstract: Some of the most astonishing and prominent properties of Quantum Mechanics, such as entanglement and Bell nonlocality, have only been studied extensively in dedicated low-energy laboratory setups. The feasibility of these studies in the high-energy regime explored by particle colliders was only recently shown, and has gathered the attention of the scientific community. For the range of particles and fundamental interactions involved, particle colliders provide a novel environment where quantum information theory can be probed, with energies exceeding, by about 12 orders of magnitude, the laboratory setups typically used in the field. Furthermore, collider detectors have inherent advantages in performing certain quantum information measurements, and allow for the reconstruction the state of the system under consideration via quantum state tomography. Here, we elaborate on the potential, challenges, and goals of this innovative and rapidly evolving line of research, and discuss its expected impact on both quantum information theory and high-energy physics.
@@ 1  Scientific context
Entanglement, the hallmark and one of the most perplexing aspects of Quantum Mechanics (QM) [1, 2], has been observed in a variety of systems ranging from photon pairs to macroscopic objects [3, 4, 5, 6, 7, 8, 9, 10, 11]. In the presence of entanglement, the full description of a composite system cannot be derived from the individual descriptions of its subsystems because of quantum correlations that interconnect them. These correlations may even survive once the subsystems have been spatially separated. In the presence of these correlations, it may also happen that the probabilities of obtaining certain outcomes from independent measurements performed on the composing subsystems no longer follow the known factorization rule. This phenomenon, known as Bell nonlocality, is quantified by a Bell inequality [12], which is violated in nonlocal theories. In stark contrast, this inequality is satisfied in any local realistic theory that seeks to remedy the quirkiness of QM through the introduction of hidden variables. Bell nonlocality is central to Quantum Information Theory (QIT) [13] and its significance was highlighted with the 2022 Nobel Prize in Physics, awarded for experimental tests with entangled photons that proved that Bell inequalities are violated in nature [14, 3, 15, 16].
The possibility of conducting tests of Bell-type correlations at colliders was first explored in Refs. [17, 18, 19, 20], whereas, more recently, entanglement has been measured in top-antitop-quark pair ($t\bar{t}$) spin correlations by the ATLAS and the CMS collaborations at the Large Hadron Collider (LHC) [21, 22, 23]. Bell nonlocal states have also been observed in data pertaining to spin correlations in charmonium and $B$-meson decays [24, 25, 26]. Furthermore, Bell nonlocality has been measured in flavor oscillations of neutral $B$-meson pairs by the Belle collaboration [7]. Remarkably, until recently, quantum correlations such as entanglement and Bell nonlocality have not been extensively investigated in the high-energy regime explored in collider experiments like the LHC.
@@
The recent measurements of entanglement in $t\bar{t}$ by the ATLAS and the CMS collaborations at the LHC have paved the way for this line of studies, serving as a proof of concept for the feasibility of such measurements. Not only these are the highest-energy ever measurements of entanglement, they are also the first ones between quarks and the first ones to probe this feature of QM with fundamental particles and interactions, therefore holding great significance. It is important to note that there is a whole hierarchy of quantum correlations that can be studied. For instance, discord is a measure of non-classical correlations that can interconnect the components of a system even if they are not entangled; steering, instead, refers to a property of the state where a measurement performed on one subsystem can steer the quantum state of the other [36]. In Figure 1, the quantum correlations at the LHC in $pp\to t\bar{t}$ production are shown on the left panel, and Bell nonlocality in $pp\to ZZ$ on the right panel.
@@
The hierarchy of quantum correlations and their possible implications have not yet been studied extensively in the high-energy regime. The collider environment carries then a unique opportunity to complete this investigation owing also to the capability of measuring quantum correlations already with present detector technologies. For example, the large statistics naturally available in collider experiments are highly-demanding to achieve in standard setups, and are essential to measure discord and steering. Specifically quantum discord can be measured according to the original definition [37], which is extremely challenging in low-energy experiments. In addition, the steering ellipsoid, which captures an important amount of information about the system [38], can be experimentally reconstructed. Previously this was achieved only once by using $5\times 10^{4}$ detection events [39]. Furthermore, QM predicts a maximal value for the violation of Bell inequalities, the Cirel’son bound [40], which could be probed at the highest available energies. Exceeding the bound would demonstrate the existence of a more fundamental theory where nonlocal effects are stronger than in QM [13, 41, 42]. Finally, direct access to relativistic and massive particles allows the investigation of quantum systems formed by qubits (such as fermions), and qutrits (such as massive gauge bosons), and may also be valuable for investigating the relativistic properties of the spin operator, a fundamental yet unresolved question in QM and in QIT itself [43, 44, 45, 46, 47, 48, 49].
@@
Beyond the fundamental interest in testing quantum correlations at high energies, exploring these concepts could bring several benefits to High-Energy Physics (HEP). Notably, new and promising techniques targeting physics beyond the Standard Model (SM) have already emerged [50, 51, 52, 53, 54, 55]. Additionally, QIT methods can be utilized to explore non-perturbative aspects of Quantum Chromodynamics (QCD) [56, 57, 58, 59, 60, 61, 62, 63]. Furthermore, this research has sparked important discussions and advancements in modeling SM processes, for instance concerning the production of $t\bar{t}$ systems used to measure quantum correlations: exploring these concepts has provided new insights into non-relativistic QCD effects, such as toponium formation [64, 65, 66, 67]. Accurate modeling of these effects is also crucial for achieving precision measurements within the SM and, without leveraging variables sensitive to the entanglement between the spins of the top quarks, such effects would remain inaccessible given current reconstruction resolution limits. The Higgs boson also plays a crucial role in studying quantum correlations, serving as a source of maximally entangled qubits or qutrits that can be analyzed using QIT methods. Moreover, these studies offer additional tools for exploring Higgs couplings and $CP$ structure. [68, 69]
@@ 2  Objectives
    At collider experiments, the presence of entanglement can be made distinguishable via a procedure dubbed quantum state tomography, which allows for the full reconstruction of the quantum state of the system under investigation. Whereas entanglement is relatively understood and well quantified for bipartite systems comprising components of low dimensionality [71], a general characterization is presently missing. In relation to that, having access to relativistic multipartite states is crucial to test the relativistic regime of QIT, and it could thus significantly improve our understanding of entanglement by directly probing its nature and properties in this regime. Within HEP, this requirement translates into the necessity of final states characterized by a large number of particles generated by a common process, achievable only at a collider able to reach large collision energies [72].
[…]
@@
[…]
@@ 3  Methodology
Currently, quantum observables are measured in colliders as spin correlations between unstable particles. When a particle decays sufficiently quickly, its spin information is transferred to its decay products. This way, by measuring angular distributions between the decay products in the rest frames of the parent particles, the correlations between the original particles can be deduced, and the full spin density matrix can be reconstructed. Such measurements require precise reconstruction of the system to accurately characterize quantum effects. Additionally, it is beneficial to isolate regions of the phase space of the decaying particles to enhance the quantum correlations in the measured density matrix, which often creates more challenges.
Having reconstructed the density matrix, we can study its QIT properties, such as entanglement. Oftentimes, however, quantities can be measured more directly without needing to perform full quantum state tomography first. This is the case for entanglement, which, near production threshold for $t\bar{t}$, is given by the angular difference between the lepton, measured in the antitop rest frame, and the antilepton, measured in the top rest frame. From this angular distribution, the parameter $D$ can be extracted [27] where $D<-1/3$ indicates an entangled state.
Both the ATLAS and CMS collaborations have proven they are up to the task of detailed spin correlation measurements and have measured entanglement in the top-pair final state near the production threshold [21, 22], shown in Figure 2. The CMS collaboration has also measured entanglement in the lepton and jets final state of the top-pair at high $m_{t\bar{t}}$, and also the polarizations and spin correlations of the $t\bar{t}$ system [23]. This result is the tip of the iceberg in terms of QIT observables that can be measured in the near future. Using spin correlations to perform quantum tomography, we can expect experimental tests of quantum discord [36, 99] in top-pairs and of entanglement and Bell nonlocality in $\tau^{+}\tau^{-}$ events at the LHC [51].
Figure 2: Measurement of the entanglement marker $D$, where $D<-1/3$ indicates entanglement. Left panel: ATLAS particle-level $D$ measurement compared with various MC models. Error bars represent all uncertainties included. The entanglement limit shown in the low $m_{t\bar{t}}$ region is a conversion from its parton-level value of $D=-1/3$ to the corresponding value at particle-level [21]. Right panel: CMS parton-level $D$ measurement either including (black filled point) or not including (black open point) contribution from toponium, compared to MC predictions with (solid line) or without (dashed line) the inclusion of toponium. Inner error bars represent the statistical uncertainty, while the outer error bars represent the total uncertainty for data [22].
@@
Flavor oscillations allow for the quantum number of flavor, rather than spin, to form the building blocks of quantum states produced at colliders. These oscillations occur between a particle and its antiparticle. When a decay occurs, the decay products will identify whether it was the particle or the antiparticle at the time of decay. The reconstruction of the decay products can identify whether it was the particle or the antiparticle which decayed. Measuring this as a function of the time of the decays allows for the reconstruction of parts of the density matrix. In particular, Belle II [102] provides a high-rate production of entangled $B$-meson pairs, offering unique opportunities for entanglement studies that remain largely unexplored. Current analyses assume perfect flavor entanglement of $B$-meson pairs, with potential decoherence effects from environmental interactions or new physics mostly untested [103]. A previous Belle study provided a quantitative test of Bell nonlocality in flavor physics [7]. Similar measurements are possible also in the $D^{0}\bar{D}^{0}$ and $K^{0}\bar{K}^{0}$ systems [104].
@@ 4  Readiness and expected challenges
Another limiting factor identified in these initial measurements is the accuracy of Monte Carlo (MC) simulations, and with this, the need to carry out an extended theoretical program to control QCD and related uncertainties. The analysis strategies employed rely on the ability of MC simulations to accurately predict detector-level distributions of angular quantities, which serve as inputs for the extraction of the spin density matrix. Given the sensitivity of these measurements to all aspects of reconstructed objects, it is imperative that every step of the simulation pipeline is highly precise. This includes not only the matrix element calculations but also parton-shower modeling and hadronization. An example of how these aspects affect the simulation in top-quark pair final states of the angular distribution used to extract the quantum entanglement marker $D$ is shown in Figure 3 (left panel).
With respect to matrix element calculations, early investigations into the impact of higher-order corrections on the extraction of the spin density matrix and related observables have demonstrated their relevance, both in the extracted quantities and their potential to constrain new physics [112, 52], as shown in Figure 3 (right panel). However, such corrections have not yet been investigated in most phenomenological analyses, although several MC event generators already support at least next-to-leading-order QCD corrections to processes of interests. Further development is required to include higher-order electroweak corrections in the MC simulations. Another related challenge that must be addressed in the coming years is the change in the intrinsic structure of spin density matrices in the presence of higher-order corrections. Additionally, the simulation of parton-shower effects and the corresponding matching to fixed-order predictions are undergoing continuous refinement. Assessing the impact of parton-shower modeling on quantum-information observables and contributing to its development will be a crucial aspect of future research, with a special focus on the inclusion of exact spin correlations.
@@
As outlined in Section 1, the application of quantum information techniques to the study of quantum observables is evolving rapidly. This progress includes both theoretical advancements and experimental improvements, where continuous refinements are expected in the coming years.
@@ 5  Timeline
  • HL-LHC.
    The large statistics expected by the HL-LHC are extremely important to be able to observe rare phenomena. This is true in general, but even more to measure quantum correlations, which are typically maximal in extreme parts of phase space, or for processes with a low cross section. In particular, observing Bell nonlocal states in $t\bar{t}$ is extremely challenging [122, 123, 124, 125], since in $t\bar{t}$ it is only present after imposing extreme cuts on the invariant mass of the top-quark pairs. In the left panel of Figure 4 an example for the expected sensitivity of measuring Bell nonlocality in $t\bar{t}$ events is shown, reaching almost $5\sigma$ with the HL-LHC expected data.
@@
    Within the timescale of HL-LHC, additional measurements are expected by Belle II, a next-generation flavor physics experiment that already began collecting electron-positron collision data [102], with strong European participation. The large dataset expected to be collected will contain billions of decays of bottom mesons, charm hadrons, and $\tau$-leptons. Belle II’s improved capabilities, such as independent measurement of $B$-meson production times due to the nano-beam scheme, enable more precise tests of flavor entanglement decoherence. Additionally, Belle II can study spin correlations in $\tau$-lepton pairs and could confirm Bell nonlocality [126].
@@
    Currently there are several projects discussing the construction of a future lepton collider. Some prominent examples are: FCC-$ee$ [76], LEP III [77], CEPC [78], Muon Colliders [74], ILC [79], CLIC [80] and C3 [81]. The distinctive processes, the precise knowledge of the center-of-mass energy of the process, and the large statistics expected at lepton colliders offer many opportunities to the field. In particular, the background for processes like $e^{+}e^{-}\to Z/\gamma^{*}\to\tau^{+}\tau^{-}$ [94] and $e^{+}e^{-}\to ZH(\to\tau^{+}\tau^{-})$ [69] is expected to be small and the resolution to reconstruct the whole final state is expected to be very accurate, allowing precise measurements of quantum correlations such as entanglement, steerability and Bell nonlocality. In the right panel of Figure 4 an example for the required experimental precision to measure Bell nonlocality in $e^{+}e^{-}\to t\bar{t}$ events with $5\sigma$, as a function of the center-of-mass energy, is shown. In the left panel of Figure 5 the regions of phase space in which the quantum state of the system is Bell nonlocal in $e^{+}e^{-}\to\tau^{+}\tau^{-}$ events, as a function of the center-of-mass energy, are shown. Another remarkable example is the possibility of measuring, in a spin-entangled pair of particles, that the entanglement can increase after one particle decays [127]. This phenomenon has no analogue for stable particles, and can be measured with $t\bar{t}$ at lepton colliders with polarized beams. The increased precision that will be reached on QIT related measurements at lepton colliders implies also an enlarged sensitivity to investigate new physics effects [128], an example is shown in the right panel of Figure 5.
@@
[…]

### v3

@@
Abstract: Some of the most astonishing and prominent properties of Quantum Mechanics, such as entanglement and Bell nonlocality, have only been studied extensively in dedicated low-energy laboratory setups. The feasibility of these studies in the high-energy regime explored by particle colliders was only recently shown and has gathered the attention of the scientific community. For the range of particles and fundamental interactions involved, particle colliders provide a novel environment where quantum information theory can be probed, with energies exceeding by about 12 orders of magnitude those employed in dedicated laboratory setups. Furthermore, collider detectors have inherent advantages in performing certain quantum information measurements, and allow for the reconstruction of the state of the system under consideration via quantum state tomography. Here, we elaborate on the potential, challenges, and goals of this innovative and rapidly evolving line of research and discuss its expected impact on both quantum information theory and high-energy physics.
@@ 1 Scientific context
Entanglement, the hallmark and one of the most perplexing aspects of Quantum Mechanics (QM) [1, 2], has been observed in a variety of systems ranging from photon pairs to macroscopic objects [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. In the presence of entanglement, the full description of a composite system cannot be derived from the individual descriptions of its subsystems because of quantum correlations that interconnect them. These correlations may even survive once the subsystems have been spatially separated. In the presence of these correlations, it may also happen that the probabilities of obtaining certain outcomes from independent measurements performed on the composing subsystems no longer follow the known factorization rule. This phenomenon, known as Bell nonlocality, is quantified by a Bell inequality [14], which is violated in nonlocal theories. In stark contrast, this inequality is satisfied in any local realistic theory that seeks to remedy the quirkiness of QM through the introduction of hidden variables. Bell nonlocality is central to Quantum Information Theory (QIT) [15], and its significance was highlighted with the 2022 Nobel Prize in Physics, awarded for experimental tests with entangled photons that demonstrated that Bell inequalities are violated [16, 3, 17, 18].
Remarkably, until recently, quantum correlations such as entanglement and Bell nonlocality have not been extensively investigated in the high-energy regime explored in proton collider experiments like the LHC. The possibility of conducting tests of Bell-type correlations at colliders was first explored in Refs. [19, 20, 21, 22, 23, 24], whereas, more recently, entanglement has been measured in top-antitop-quark pair ($t\bar{t}$) spin correlations by the ATLAS and the CMS collaborations at the Large Hadron Collider (LHC) [25, 26, 27]. Bell nonlocal states have also been observed in data pertaining to spin correlations in charmonium and $B$-meson decays [28, 29, 30]. Furthermore, Bell nonlocality has been measured in flavor oscillations of neutral $B$-meson pairs by the Belle collaboration [7]. In addition, the ALICE collaboration recently measured quantum interference effect at the femtometer scale [31].
@@
The recent measurements of entanglement in $t\bar{t}$ by the ATLAS and the CMS collaborations at the LHC have paved the way for this line of studies, serving as a proof of concept for the feasibility of such measurements. Not only these are the highest-energy ever measurements of entanglement, they are also the first ones between quarks and the first ones to probe this feature of QM by utilizing unique fundamental particles and interactions, therefore holding great significance. It is important to note that there is a whole hierarchy of quantum correlations that can be studied. For instance, discord is a measure of non-classical correlations that can interconnect the components of a system even if they are not entangled; steering, instead, refers to a property of the state where a measurement performed on one subsystem can steer the quantum state of the other [41]. In Fig. 1, the quantum correlations at the LHC in $pp\to t\bar{t}$ production are shown on the left panel, and Bell nonlocality in $pp\to ZZ$ on the right panel.
@@
The hierarchy of quantum correlations and their possible implications have not yet been studied extensively in the high-energy regime. The collider environment carries then a unique opportunity to complete this investigation owing also to the capability of measuring quantum correlations already with present detector technologies. For example, measurements such as quantum discord and steering require large data samples. These are already accessible in specific processes, such as top-quark pair final states, within the datasets already collected at the LHC and those expected in upcoming runs. In contrast, achieving comparable statistics in standard experimental setups remains challenging. Specifically, quantum discord can be measured according to the original definition [42], which is extremely challenging in low-energy experiments. In addition, the steering ellipsoid, which captures an important amount of information about the system [43], can be experimentally reconstructed. Previously this was achieved only once by using $5\times 10^{4}$ detection events [44]. Furthermore, QM predicts a maximal value for the violation of Bell inequalities, the Cirel’son bound [45], which could be probed at the highest available energies. Exceeding the bound would demonstrate the existence of a more fundamental theory where nonlocal effects are stronger than in QM [15, 46, 47]. Finally, direct access to relativistic and massive particles allows the investigation of quantum systems formed by qubits (such as fermions), and qutrits (such as massive gauge bosons), and may also be valuable for investigating the relativistic properties of the spin operator, a fundamental yet unresolved question in QM and in QIT itself [48, 49, 50, 51, 52, 53, 54].
@@
Beyond the fundamental interest in testing quantum correlations at high energies, exploring these concepts could bring several benefits to High-Energy Physics (HEP). Notably, new and promising techniques targeting physics beyond the Standard Model (SM) have already emerged [55, 56, 57, 58, 59, 60]. Additionally, QIT methods can be utilized to explore non-perturbative aspects of Quantum Chromodynamics (QCD) [61, 62, 63, 64, 65, 66, 67, 68]. Furthermore, this research has sparked important discussions and advancements in modeling SM processes, for instance, concerning the production of $t\bar{t}$ systems used to measure quantum correlations: exploring these concepts has provided new insights into non-relativistic QCD effects, such as toponium formation [69, 70, 71, 72]. Accurate modeling of these effects is also crucial for achieving precision measurements within the SM and, without leveraging variables sensitive to the entanglement between the spins of the top quarks, such effects would remain inaccessible given current reconstruction resolution limits. The Higgs boson also plays a crucial role in studying quantum correlations, serving as a source of maximally entangled qubits or qutrits that can be analyzed using QIT methods. Moreover, these studies offer additional tools for exploring Higgs couplings and $CP$ structure. [73, 74]
@@ 2 Objectives
    At collider experiments, the presence of entanglement can be made distinguishable via quantum state tomography, which allows for the full reconstruction of the quantum state of the system under investigation. Whereas entanglement is relatively understood and well quantified for bipartite systems comprising components of low dimensionality [76], a general characterization is presently missing. In relation to that, having access to relativistic multipartite states is crucial to test the relativistic regime of QIT, and it could thus significantly improve our understanding of entanglement by directly probing its nature and properties in this regime. Within HEP, this requirement translates into the necessity of final states characterized by a large number of particles generated by a common process, achievable only at a collider able to reach large collision energies [77].
    In light of this, FCC-$hh$ [78], a multi-TeV muon collider [79], and the SPPC [80] seem the ideal tools to pursue this line of studies. Precision machines such as FCC-$ee$ [81], LEP III [82], CEPC [83], or the proposed linear colliders instead [84, 85, 86], could be used to achieve extremely accurate tomographic measurements that could help to highlight any deviation from the SM predictions. On a more theoretical ground, we also remark that the literature offers several entanglement measures, monotones and witnesses apt to investigate the phenomenon [87]. Their meaning and the utility within particle physics is to be investigated and clarified.
@@
    The particle physics community has started to ascertain which processes are more suitable for this kind of analysis at present and future colliders; however, the analysis is far from being exhaustive. Presently, the ATLAS and CMS collaborations are forced to use unstable particles to investigate entanglement in spin correlations involving top quarks, bottom quarks, $\tau$-leptons, as well as $W$- and $Z$-bosons. Likewise, similar analyses can be performed with the LHCbb detector by reconstructing the decay chains initiated, for instance, by $B$-mesons. Entanglement could also be investigated at LHC experiments through the $B^{0}\bar{B}^{0}$, $D^{0}\bar{D}^{0}$, and $K^{0}\bar{K}^{0}$ flavor oscillations, which offer a complementary probe of quantum behavior [93]. Stable particles such as electrons, muons and photons could, in principle, also be used for these kinds of analyses provided that the future detector technologies are developed enough to measure their spin directly.
@@ 3 Methodology
Currently, quantum observables are measured in colliders as spin correlations between unstable particles. When a particle decays sufficiently quickly, its spin information is transferred to its decay products. In this way, by measuring angular distributions between the decay products in the rest frames of the parent particles, the correlations between the original particles can be deduced, and the full spin density matrix can be reconstructed. Such measurements require precise reconstruction of the system to accurately characterize quantum effects. Additionally, it is beneficial to isolate regions of the phase space of the decaying particles to enhance the quantum correlations in the measured density matrix, which often creates more challenges.
Having reconstructed the density matrix, we can study its QIT properties, such as entanglement. Oftentimes, however, quantities can be measured more directly without needing to perform full quantum state tomography first. This is the case for entanglement, which, near production threshold for $t\bar{t}$, is given by the angular difference between the lepton, measured in the antitop rest frame, and the antilepton, measured in the top rest frame. From this angular distribution, the parameter $D$ can be extracted [32] where D<1/3D<-1/3 indicates an entangled state.
Both the ATLAS and CMS collaborations have proven they are up to the task of detailed spin correlation measurements and have measured entanglement in the top-pair final state near the production threshold [25, 26], shown in Fig. 2. The CMS collaboration has also measured entanglement in the lepton and jets final state of the top-pair at high $m_{t\bar{t}}$, and also the polarizations and spin correlations of the $t\bar{t}$ system [27]. This result is the tip of the iceberg in terms of QIT observables that can be measured in the near future. Using spin correlations to perform quantum tomography, we can expect experimental tests of quantum discord [41, 104] in top-pairs and of entanglement and Bell nonlocality in $\tau^{+}\tau^{-}$ events at the LHC [56, 105].
Figure 2: Measurement of the entanglement marker $D$, where D<1/3D<-1/3 indicates entanglement. Left panel: ATLAS particle-level $D$ measurement compared with various MC models. Error bars represent all uncertainties included. The entanglement limit shown in the low $m_{t\bar{t}}$ region is a conversion from its parton-level value of D=1/3D=-1/3 to the corresponding value at particle-level [25]. Right panel: CMS parton-level $D$ measurement either including (black filled point) or not including (black open point) contribution from toponium, compared to MC predictions with (solid line) or without (dashed line) the inclusion of toponium. Inner error bars represent the statistical uncertainty, while the outer error bars represent the total uncertainty for data [26].
@@
Flavor oscillations allow for the quantum number of flavor, rather than spin, to form the building blocks of quantum states produced at colliders. These oscillations occur between a particle and its antiparticle. When a decay occurs, the decay products will identify whether it was the particle or the antiparticle at the time of decay. The reconstruction of the decay products can identify whether it was the particle or the antiparticle which decayed. Measuring this as a function of the time of the decays allows for the reconstruction of parts of the density matrix. In particular, Belle II [108] provides a high-rate production of entangled $B$-meson pairs, offering unique opportunities for entanglement studies that remain largely unexplored. Current analyses assume perfect flavor entanglement of $B$-meson pairs, with potential decoherence effects from environmental interactions or new physics mostly untested [109]. A previous Belle study provided a quantitative test of Bell nonlocality in flavor physics [7]. Similar measurements are possible also in the $D^{0}\bar{D}^{0}$ and $K^{0}\bar{K}^{0}$ systems [110, 13, 24] and novel effects are under study [111].
@@ 4 Readiness and expected challenges
Another limiting factor identified in these initial measurements is the accuracy of Monte Carlo (MC) simulations, and with this, the need to carry out an extended theoretical program to control QCD and related uncertainties. The analysis strategies employed rely on the ability of MC simulations to accurately predict detector-level distributions of angular quantities, which serve as inputs for the extraction of the spin density matrix. Given the sensitivity of these measurements to all aspects of reconstructed objects, it is imperative that every step of the simulation pipeline is highly precise. This includes not only the matrix element calculations but also parton-shower modeling and hadronization. An example of how these aspects affect the simulation in top-quark pair final states of the angular distribution used to extract the quantum entanglement marker $D$ is shown in Fig. 3 (left panel).
With respect to matrix element calculations, early investigations into the impact of higher-order corrections on the extraction of the spin density matrix and related observables have demonstrated their relevance, both in the extracted quantities and their potential to constrain new physics [119, 120, 57], as shown in Fig 3 (right panel). However, such corrections have not yet been investigated in most phenomenological analyses, although several MC event generators already support at least next-to-leading-order QCD corrections to processes of interests. Further development is required to include higher-order electroweak corrections in the MC simulations. Another related challenge that must be addressed in the coming years is the change in the intrinsic structure of spin density matrices in the presence of higher-order corrections. Additionally, the simulation of parton-shower effects and the corresponding matching to fixed-order predictions are undergoing continuous refinement. Assessing the impact of parton-shower modeling on quantum-information observables and contributing to its development will be a crucial aspect of future research, with a special focus on the inclusion of exact spin correlations.
@@
As outlined in Sect. 1, the application of quantum information techniques to the study of quantum observables is evolving rapidly. This progress includes both theoretical advancements and experimental improvements, where continuous refinements are expected in the coming years.
@@ 5 Timeline
  • HL-LHC and Belle II.
    The large statistics expected by the HL-LHC are extremely important to be able to observe rare phenomena. This is true in general, but even more to measure quantum correlations, which are typically maximal in extreme parts of phase space, or for processes with a low cross section. In particular, observing Bell nonlocal states in $t\bar{t}$ is extremely challenging [130, 131, 132, 133], since in $t\bar{t}$ it is only present after imposing extreme cuts on the invariant mass of the top-quark pairs. In the left panel of Fig. 4 an example for the expected sensitivity of measuring Bell nonlocality in $t\bar{t}$ events is shown, reaching almost $5\sigma$ with the HL-LHC expected data.
@@
    Within the timescale of HL-LHC, additional measurements are expected by Belle II, a next-generation flavor physics experiment that already began collecting electron-positron collision data [108], with strong European participation [134]. The large dataset expected to be collected will contain billions of decays of bottom mesons, charm hadrons, and $\tau$-leptons. Belle II’s improved capabilities, such as independent measurement of $B$-meson production times due to the nano-beam scheme, enable more precise tests of flavor entanglement decoherence. Additionally, Belle II can study spin correlations in $\tau$-lepton pairs and could confirm Bell nonlocality [135].
@@
    Currently there are several projects discussing the construction of a future lepton collider. Some prominent examples are: FCC-$ee$ [81], LEP III [82], CEPC [83], Muon Colliders [79], ILC [84], CLIC [85] and C3 [86]. The distinctive processes, the precise knowledge of the center-of-mass energy of the process, and the large statistics expected at lepton colliders offer many opportunities to the field. In particular, the background for processes like $e^{+}e^{-}\to Z/\gamma^{*}\to\tau^{+}\tau^{-}$ [99], $e^{+}e^{-}\to ZH(\to\tau^{+}\tau^{-})$ [74], $e^{+}e^{-}\to ZH(\to ZZ^{*})$ [136], $\mu^{+}\mu^{-}\to ZZ$ [137] and $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}H(\to ZZ^{*})$ [138] is expected to be small and the resolution to reconstruct the whole final state is expected to be very accurate, allowing precise measurements of quantum correlations such as entanglement, steerability and Bell nonlocality. In the right panel of Fig. 4 an example for the required experimental precision to measure Bell nonlocality in $e^{+}e^{-}\to t\bar{t}$ events with $5\sigma$, as a function of the center-of-mass energy, is shown. In the left panel of Fig. 5 the regions of phase space in which the quantum state of the system is Bell nonlocal in $e^{+}e^{-}\to\tau^{+}\tau^{-}$ events, as a function of the center-of-mass energy, are shown. Another remarkable example is the possibility of measuring, in a spin-entangled pair of particles, that the entanglement can increase after one particle decays [139]. This phenomenon has no analogue for stable particles, and can be measured with $t\bar{t}$ at lepton colliders with polarized beams. The increased precision that will be reached on QIT related measurements at lepton colliders implies also an enlarged sensitivity to investigate new physics effects [140], an example is shown in the right panel of Fig. 5.
@@
    So far, most of the theoretical studies focused on bipartite systems. One of the main reasons for this is that at the LHC, and at future lepton colliders, the cross section for the production of more than two heavy particles that decay quickly enough to propagate their spin information to their decays products, is small. Nevertheless, in new machines targeting a center-of-mass energy of the order of a hundred TeV, such as FCC-$hh$ [78] or SPPC [80], these processes will be more abundantly produced. This presents an opportunity to study entanglement in multi-particle systems, a challenging but also very interesting possibility [77, 141]. However, more work on the theory side is required in order to better understand and measure relativistic multi-particle quantum correlations.